OpenStudy (anonymous):

Solve: 11x-6y^2+7y= -14 x= -7y + 3

OpenStudy (anonymous):

put the equation for x into the top equation and solve

mathslover (mathslover):

\[11(-7y+3)-6y^2+7y=-14\] \[-77y+33-6y^2+7y=-14\] \[-70y-6y^2=-14-33\] \[-70y-6y^2=-47\] \[6y^2+70y-47=0\] solve this quadratic equation

OpenStudy (precal):

use the quadratic formula, it would be easier and faster

mathslover (mathslover):

right @precal

OpenStudy (anonymous):

is it -35+ radical 1507 over 6 ?

OpenStudy (anonymous):